# Recov-Vision: Linking Street View Imagery and Vision-Language Models for Post-Disaster Recovery

### v1

@@ III-B Data pre-processing
III-B1 View rectification
Figure 2: Visualization of the view rectification process. (a) The top-down view illustrates the geometric parameters used for rectification, including the vehicle’s position $(x_{1},y_{1})$ and heading $(\theta_{1})$, and the building’s centroid $(x_{2},y_{2})$. The bearing to the building $(\theta_{2})$ is calculated, and the yaw angle $\alpha$ orients the rectified field of view (green) toward the building. (b) A raw 360-degree panoramic image is processed to generate a rectified, planar view of the building facade, which is then used for analysis.
@@ III-B1 View rectification
$x$ $=\cos(lat_{1})\cdot\sin(lat_{2})$
$\quad-\sin(lat_{1})\cdot\cos(lat_{2})\cdot\cos(\Delta lon)$ (3)
$\theta$ $=\text{atan2}(y,x)\cdot\frac{180}{\pi}$ (4)
@@
The resulting rectified facade views constitute the inputs to the VLM prompting module (Section III-C) and the temporal comparison and recovery mapping (Section III-D). Calibration offset, field-of-view, and window size are tunable; reported analyses use values set a priori for the study area and collection period.
@@
III-C VLM prompting strategy
The proposed building occupancy assessment in this work employs vision-language models to analyze the rectified facade images from the pre-processing pipeline. Two distinct strategies are implemented and compared: a single-stage vision-only approach and a two-stage vision-reasoning pipeline (Figure 3).
@@ III-C VLM prompting strategy
Thus, the effective threshold is two when no vehicle is visible and three when a vehicle is present (i.e., one point is subtracted for visible vehicles to reflect the prior that cars correlate with occupancy). Table I summarizes the rules used in all experiments.
TABLE I: Baseline threshold ($\tau$) with vehicle adjustment. Risk indicators are: house_destruction, structural_damage, exterior_debris, open_doors_windows, site_accessible=false, exterior_mud, emergency_markings, major_repairs.
@@
The second strategy introduces a separate decision-making stage. In the feature-extraction stage, the same multi-modal model analyzes each rectified image with a prompt that requests only observable visual evidence (the nine attributes above), without inferring occupancy. In the subsequent decision stage, we use the language model (GPT-4o) in text-only mode to map the extracted attributes to a final occupancy label via explicit reasoning rules and few-shot learning examples (in the Appendix section). These rules and exemplars (nine cases covering diverse damage and occupancy scenarios) encode domain expertise. The decision policy is intentionally conservative, favoring “Not Occupied” when evidence is ambiguous or suggests potential habitability concerns. This two-stage approach is referred to as the “Two-stage” strategy in the following analysis.
This two-stage approach offers several methodological advantages over direct classification. The separation of perception and reasoning enables systematic evaluation of visual feature extraction accuracy independent of decision logic. The structured attribute representation provides interpretable intermediate outputs that facilitate error analysis and model debugging. The explicit reasoning stage allows incorporation of domain knowledge and conservative decision policies without requiring retraining of the vision components. Additionally, the standardized attribute schema enables consistent evaluation across different building types and damage patterns. The attributes used in the VLM extraction stage and their typical visual cues are summarized in Table in the Appendix section. Note that Emergency responders often spray-paint a large “X” on building exteriors during search-and-rescue operations, indicating that the structure was inspected and often that it was damaged or uninhabitable.
@@
All language-model calls used the OpenAI API via a chat/responses endpoint that accepts role-based messages and images. For fairness, both strategies shared an identical vision–attribute extraction stage (same multimodal model, standard schema, and prompt); only the decision stages are different. The vision extractor was invoked with a system message and a user message (image + standard schema) at temperature=0, max_tokens=200; the decision stage used a text-only call at temperature=0, max_tokens=5. No explicit response_format (JSON mode) was requested—JSON structure was enforced by instruction. The one-stage baseline applied a deterministic scoring rule to the extracted attributes. The two-stage method used a text-only LLM decision with the system message “You are a helpful assistant that analyzes images and outputs structured JSON.”, $\texttt{temperature}=0$, max_tokens=5, and the few-shot exemplars in the Appendix section. No seed was set, and other sampling parameters used defaults (e.g., top_p). The code does not implement automatic JSON validation or retry-on-failure; responses were consumed as returned.
III-D Temporal comparison and recovery mapping
This stage derives parcel‑level occupancy change descriptors between the two field campaigns for each inference strategy and prepares data structures for recovery mapping. Inputs are per‑parcel visit labels from the VLM prompting stage (Section III-C) for both the single‑stage (vision scoring) and two‑stage (vision + reasoning) strategies, together with human ground‑truth annotations for the subset of parcels manually labeled in each campaign. After visibility and quality filtering, parcel identifiers are joined across visits in chronological order (V1 for Visit 1 and V2 for Visit 2). When multiple rectified frames exist for a parcel within a visit, each strategy first produces frame‑level labels; a visit‑level label is then assigned by majority vote. Ties, mutually contradictory evidence, or any frame flagged Uncertain produce a conservative Not Occupied assignment; if all usable frames are Uncertain, the visit label is set to Uncertain and excluded from primary accuracy and change metrics. The same consolidation rule is applied to ground‑truth frame annotations.
For parcels with valid (Occupied or Not Occupied) labels at both visits, we define a change class given by the ordered pair (V1, V2). The four resulting classes are: Recovered (from Not Occupied to Occupied), Deteriorated (from Occupied to Not Occupied), Stable–Occupied (stay as Occupied), and Stable–Not Occupied (stay as Not Occupied), summarized in Table II. Parcels with an Uncertain or missing label at either visit are excluded from change summaries for that strategy but still counted in coverage statistics (fraction of parcels yielding a usable change).
TABLE II: Parcel occupancy change classes between Visit 1 (V1) and Visit 2 (V2).
@@ III-D Temporal comparison and recovery mapping
Structured outputs of this stage are: (1) a visit‑level label table containing, for each parcel and strategy (and ground truth where available), the consolidated V1 and V2 labels with exclusion flags; and (2) a parcel change table enumerating change classes for parcels with valid labels at both visits. These outputs serve as inputs to the evaluation metrics and numerical results in Section IV.
III-E Model Performance
@@
IV Results
IV-A VLM prompting pipeline performance
The original dataset included 578 building facades from Visit 1 and 621 from Visit 2. After filtering out images labeled as “Unknown” due to obstruction, blur, poor viewing angle, or no building present, the usable sample comprised 471 facades for Visit 1 and 443 for Visit 2. These filtered counts define the set of parcels used for model evaluation across both strategies. Model performance is compared to ground truth in several ways. The Ground Truth, predictions by the Two-stage strategy, and by the One-stage strategy are shown in Figure 4. The figure shows the occupancy status for each parcel at both visits. Ground truth vacancies declined from 85 to 77 (–8; –9.4%), indicating a modest decrease in vacancies between the two field surveys. The Two-stage strategy shows the same trend but with greater magnitude (75 to 63; –12; –16.0%), while the One-stage strategy shows an even larger drop (70 to 54; –16; –22.9%). However, both models overestimate occupancy relative to ground truth at each visit (Two-stage: –10 V1 / –14 V2; One-stage: –15 V1 / –23 V2), with larger overestimates at V2 consistent with under-predicted vacancy. Total counted parcels differ between visits because parcels with Uncertain or missing labels are excluded, but the relative shifts still show an occupancy contraction overstated by the One-stage baseline. These distributional biases signal the net recovery bias and change agreement results reported below.
@@ IV-A VLM prompting pipeline performance
For each strategy, we compute per‑visit confusion matrices (shown in Figure 5; Not Occupied is treated as positive class) and report accuracy, precision, recall, F1, and Cohen’s $\kappa$ in Table IV. Table IV summarizes performance with 95% CIs, and Table III reports paired significance tests.
TABLE III: Statistical significance tests for method comparisons. McNemar tests compare per-visit error patterns; paired bootstrap tests compare metric differences.
@@
TABLE IV: Method performance comparison with bootstrap confidence intervals. Not Occupied is treated as the positive class.
@@
As shown in Table IV, the two-stage strategy exhibits a modest precision-recall trade-off relative to the one-stage baseline. It achieves higher recall at the cost of slightly lower precision, resulting in numerically higher F1 and Cohen’s $\kappa$ scores. However, as shown in Table III, these differences are not statistically significant. Both McNemar’s tests on per-visit error patterns and paired bootstrap tests on the performance metrics yield non-significant p-values, with confidence intervals for the differences spanning zero. This suggests that while the two-stage method yields slightly higher recall and agreement, the two strategies are statistically indistinguishable at the current operating point.
Figure 6: Per‑visit overall accuracy: (a) One-stage strategy (95.5% V1, 92.1% V2); (b) Two-stage strategy (96.2% V1, 93.2% V2).
Overall accuracies (Figure 6) decline modestly from V1 to V2 for both strategies (Two-stage -3.0 pp (percentage point) vs One-stage -3.4 pp), with the Two-stage maintaining a consistent  0.7–1.1 pp advantage each visit. The parallel drop suggests a shared increase in classification difficulty (e.g., more subtle post‑repair cues) rather than strategy‑specific drift.
Direct comparison of the two strategies focuses on expected advantages of the two‑stage (vision + reasoning) method: fewer false Not Occupied labels at a single visit (improved Occupied recall), higher precision for Deteriorated changes by propagating fewer spurious vacancies, improved precision for Recovered (fewer artificial recoveries caused by an incorrect Not Occupied baseline at V1), and higher overall change $\kappa$. Disagreements between strategies are flagged for manual review and spatially highlighted.
@@
IV-B Occupancy change analysis
We analyze parcel‑level occupancy change between V1 and V2 using the four classes defined in Table II (Stable Occupied, Stable Not Occupied, Recovered, Deteriorated). Figure 9 (a)–(c) show Recovered / Deteriorated counts: ground truth 20 / 6 (net +14), Two-stage 23 / 9 (net +14; matches net recovery), One-stage 24 / 6 (net +18; overstates recovery by +4). Thus, both strategies capture the direction of net improvement; the One-stage inflates recoveries, while the Two-stage balances additional true recoveries with extra deteriorations.
@@ IV-B Occupancy change analysis
Figure 9: Change analysis. (a) Ground‑truth Recovered vs. Deteriorated counts (n=26). (b) One-stage predicted changes (n=30). (c) Two-stage predicted changes (n=32).
TABLE V: Parcel-level occupancy change agreement categories (confident parcels with both visits).
@@
GT + Two-stage Agree 8 3.3
GT + One-stage Agree 4 1.7
@@
Spatial heterogeneity and remaining uncertainty are visualized in Figure 10, which maps (a) Visit 1 ground‑truth occupancy, (b) Visit 2 ground‑truth occupancy, (c) ground‑truth change classes (Recovered, Deteriorated, Unchanged), and (d) spatial distribution of Two-stage prediction accuracy (both visits correct, one visit correct, both wrong). This multi‑panel view highlights localized pockets of recovery and deterioration and reveals where residual model errors cluster geographically, indicating priority areas for data or prompt refinement.
Figure 10: Geographic distribution of occupancy and model accuracy: (a) Visit 1 ground‑truth status; (b) Visit 2 ground‑truth status; (c) ground‑truth change classes (Recovered, Deteriorated, Unchanged); (d) Two-stage prediction accuracy by parcel (both visits correct, one visit correct, both wrong).
@@
We assessed spatial autocorrelation with the global Moran’s $I$ for four spatial fields: parcel occupancy at Visit 1 and Visit 2, the ground-truth change field, and a binary accuracy indicator (1 = correct, 0 = incorrect). Positive $I$ with small $p$-values indicates spatial clustering rather than spatial randomness. Table VI summarizes the results.
TABLE VI: Global Moran’s II statistics for parcel-level layers. Positive values indicate spatial clustering. The expected value under the randomization null is E[I]=1/(n1)E[I]=-1/(n-1); here $E[I]\approx-0.0042$ because $n=240$.
@@
Visit 1 occupancy 0.3704 0.0042-0.0042 0.0000
Visit 2 occupancy 0.2834 0.0042-0.0042 0.0000
@@
V Discussion
Aggregate comparisons with uncertainty and paired tests indicate that, at our chosen operating point, the two-stage strategy achieves higher recall and slightly higher F1 and Cohen’s $\kappa$ than the one-stage baseline, at the cost of a small reduction in precision. These differences are not statistically significant (McNemar and paired bootstrap), so we regard the two strategies as comparable in overall agreement. The operational value of the two-stage design lies in interpretability and controllability: exposing intermediate reasoning and attribute signals supports risk-aware thresholding and targeted review of spatial error clusters. While our method measures parcel-level occupancy change, these shifts provide a practical proxy for community recovery when interpreted alongside other indicators.
Beyond accuracy, the approach reduces survey time and labor. Across two campaigns we produced usable facades for 914 residential parcels at roughly 90 parcels per survey hour. It is an order-of-magnitude gain over house-to-house assessments, which reinforces its value as a scalable complement to manual surveys when rapid neighborhood information is required.
Accuracy declined from Visit 1 to Visit 2 for both methods, likely due to subtler post-repair cues and weaker signals. Few parcels were uniquely misclassified by the reasoning approach, suggesting the added layer increased signal rather than noise. Extending to longer-term recovery monitoring will require richer temporal context beyond facade imagery alone.
Change detection highlights the benefits of reasoning. Both approaches captured the direction of occupancy, but the baseline exaggerated improvement by propagating an inflated vacancy estimate. The reasoning strategy balanced additional true recoveries with a few deteriorations, producing a more faithful estimate of net change—important for field tasking and resource planning.
Spatial mapping adds interpretability by revealing residual errors clustered in small pockets, while large areas show consistent agreement with ground truth. These patterns likely reflect local building or damage typologies and enable efficient triage by directing review to high‑uncertainty regions. The pipeline scales raw street‑view video into parcel‑level states, limiting human oversight to ambiguous cases, and its modular prompts and model weights support adaptation as models improve.
Limitations include evaluation in a single event area and exclusion of uncertain or missing labels. Only two visits constrain temporal inference and may miss nonlinear recovery (e.g., temporary re‑vacancy), and environmental variation (illumination, foliage, seasonality) was not modeled, limiting transferability over time. Reasoning traces were not systematically audited, and both strategies share the same VLM backbone, so results reflect prompt and architectural choices rather than new model training.
These limitations suggest next steps: (i) additional visits and simple temporal models; (ii) using spatial disagreement maps to target data collection; (iii) prompt and attribute refinements for recurring failure modes (e.g., boarded or tinted windows); (iv) fusing VLM outputs with GIS/building attributes and calibrated probabilities for risk-aware thresholds; and (v) auditing reasoning traces and codifying privacy policies.
In summary, inserting a concise reasoning stage yields more balanced and directionally accurate monitoring of post-disaster vacancy than a single-stage baseline. Improvements concentrate on vacancy recall and net-recovery estimation, while residual errors remain geographically localized and thus tractable for human review. Expanding geographic scope, adding temporal depth and uncertainty, and stress-testing environmental robustness are key steps toward operational deployment.
VI Conclusion
Automated occupancy monitoring is a crucial step toward rapid recovery tracking. Leveraging street-view imagery (SVI) makes this operational because it observes the facade and access cues—blocked entries, mud lines, emergency placards, active repairs—that nadir imagery cannot see. We presented a street-level, parcel-scale pipeline that pairs panoramic rectification with vision–language prompting to assess post-disaster building occupancy and track change between visits, comparing a transparent one-stage rule with a two-stage design that separates perception from conservative reasoning.
Across two visits, both strategies achieved high agreement with ground truth. The two-stage method traded a slight precision drop for higher recall and reproduced the ground-truth net recovery (+14), while the one-stage baseline overstated improvement. Spatial diagnostics (global Moran’s $I$ and error maps) revealed significant clustering of both occupancy states and residual errors, enabling targeted review of localized pockets.
Operationally, separating perception and reasoning surfaces interpretable intermediate signals and supports risk-aware thresholding and human-in-the-loop oversight. Limitations include a single event area, two field visits, exclusion of uncertain labels, unmodeled environmental variation, and reliance on one VLM backbone. Next steps include expanding geographic coverage, adding temporal depth and calibrated probabilities, auditing reasoning traces, and fusing VLM outputs with lightweight GIS and building attributes.
In sum, the approach provides a scalable way to convert raw street-view video into parcel-level occupancy states and recovery indicators. While aggregate differences between strategies are not statistically significant, the two-stage design offers practical advantages in interpretability and controllability, and the spatial analyses identify where to focus limited field and labeling resources for reliable, operational deployment.

### v2

@@ 3.2 Data pre-processing
3.2.1 View rectification
(a) Geometric relationship for view rectification.
(b) Rectification of a panoramic image to a planar projection.
Figure 2: Visualization of the view rectification process. (a) The top-down view illustrates the geometric parameters used for rectification, including the vehicle’s position $(x_{1},y_{1})$ and heading $(\vec{\theta}_{1})$, and the building’s centroid $(x_{2},y_{2})$. The bearing to the building $(\vec{\theta}_{2})$ is calculated, and the yaw angle $\alpha$ orients the rectified field of view (green) toward the building. (b) A raw 360-degree panoramic image is processed to generate a rectified, planar view of the building facade, which is then used for analysis.
@@ 3.2.1 View rectification
$x$ $=\cos(lat_{1})\cdot\sin(lat_{2})-\sin(lat_{1})\cdot\cos(lat_{2})\cdot\cos(\Delta lon)$ (3)
$\theta$ $=\text{atan2}(y,x)\cdot\frac{180}{\uppi}$ (4)
@@
The resulting rectified facade views constitute the inputs to the VLM prompting module (Section 3.3) and the temporal comparison and recovery mapping (Section 3.4). Calibration offset, field-of-view, and window size are tunable; reported analyses use values set a priori for the study area and collection period.
@@
3.3 VLM prompting strategy
The proposed building occupancy assessment employs vision-language models to analyze the rectified facade images from the pre-processing pipeline. Two distinct strategies are implemented and compared: a single-stage vision-only approach and a two-stage vision-reasoning pipeline (Figure 3).
@@ 3.3 VLM prompting strategy
Thus, the effective threshold is two when no vehicle is visible and three when a vehicle is present (i.e., one point is subtracted for visible vehicles to reflect the prior that cars correlate with occupancy). Table 1 summarizes the rules used in all experiments.
Table 1: Baseline threshold ($\tau$) with vehicle adjustment. Risk indicators are: house_destruction, structural_damage, exterior_debris, open_doors_windows, site_accessible=false, exterior_mud, emergency_markings, major_repairs.
@@
The second strategy introduces a separate decision-making stage. In the feature-extraction stage, the same multi-modal model analyzes each rectified image with a prompt that requests only observable visual evidence (the nine attributes above), without inferring occupancy. In the subsequent decision stage, we use the language model (GPT-4o) in text-only mode to map the extracted attributes to a final occupancy label via explicit reasoning rules and few-shot learning examples (see Appendix A, Listing 3). These rules and exemplars (nine cases covering diverse damage and occupancy scenarios) encode domain expertise. The decision policy is intentionally conservative, favoring “Not Occupied” when evidence is ambiguous or suggests potential habitability concerns. This two-stage approach is referred to as the “Two-stage” strategy in the following analysis.
Table 2: Attribute schema used in VLM extraction: definitions and typical observable cues.
Attribute in prompts Typical visual cues (examples)
emergency_markings Official post-disaster markings: spray-painted search/inspection tags (e.g., X-codes), placards, colored notices on doors/windows; exclude ordinary numbers or graffiti.
This two-stage approach offers several methodological advantages over direct classification. The separation of perception and reasoning enables systematic evaluation of visual feature extraction accuracy independent of decision logic. The structured attribute representation provides interpretable intermediate outputs that facilitate error analysis and model debugging. The explicit reasoning stage allows incorporation of domain knowledge and conservative decision policies without requiring retraining of the vision components. Additionally, the standardized attribute schema enables consistent evaluation across different building types and damage patterns. The attributes used in the VLM extraction stage and their typical visual cues are summarized in Table 2. Note that Emergency responders often spray-paint a large “X” on building exteriors during search-and-rescue operations, indicating that the structure was inspected and often that it was damaged or uninhabitable.
@@
All language-model calls used the OpenAI API (chat/responses endpoint), which accepts role-based messages and images. For fairness, both strategies shared the same vision attribute-extraction stage (identical multimodal model, schema, and prompt; see Appendix A, Listings 12); only the decision stage differed. The vision extractor used a system message and a user message (image plus schema) with temperature=0 and max_tokens=200. The decision stage used a text-only call with temperature=0 and max_tokens=5, along with the system message “You are a helpful assistant that analyzes images and outputs structured JSON.” and the few-shot exemplars in Listing 3. No explicit response_format (JSON mode) was requested; instead, JSON structure was enforced by instruction. The one-stage baseline then applied a deterministic scoring rule to the extracted attributes.
No random seed was set, and other sampling parameters used default values (e.g., top_p). The code does not implement automatic JSON validation or retry-on-failure; model responses were consumed as returned.
3.4 Temporal comparison and recovery mapping
This stage derives parcel-level occupancy change descriptors between the two field campaigns for each inference strategy and prepares data structures for recovery mapping. Inputs are per-parcel visit labels from the VLM prompting stage (Section 3.3) for both the single-stage (vision scoring) and two-stage (vision and reasoning) strategies, together with human ground-truth annotations for the subset of parcels manually labeled in each campaign. After visibility and quality filtering, parcel identifiers are joined across visits in chronological order (V1 for Visit 1 and V2 for Visit 2). When multiple rectified frames exist for a parcel within a visit, each strategy first produces frame-level labels; a visit-level label is then assigned by majority vote. Ties, mutually contradictory evidence, or any frame flagged Uncertain produce a conservative Not Occupied assignment; if all usable frames are Uncertain, the visit label is set to Uncertain and excluded from primary accuracy and change metrics. The same consolidation rule is applied to ground-truth frame annotations.
For parcels with valid (Occupied or Not Occupied) labels at both visits, we define a change class given by the ordered pair (V1, V2). The four resulting classes are: Recovered (from Not Occupied to Occupied), Deteriorated (from Occupied to Not Occupied), Stable–Occupied (stay as Occupied), and Stable–Not Occupied (stay as Not Occupied), summarized in Table 3. Parcels with an Uncertain or missing label at either visit are excluded from change summaries for that strategy but still counted in coverage statistics (fraction of parcels yielding a usable change).
Table 3: Parcel occupancy change classes between Visit 1 (V1) and Visit 2 (V2).
@@ 3.4 Temporal comparison and recovery mapping
Structured outputs of this stage are: (1) a visit-level label table containing, for each parcel and strategy (and ground truth where available), the consolidated V1 and V2 labels with exclusion flags; and (2) a parcel change table enumerating change classes for parcels with valid labels at both visits. These outputs serve as inputs to the evaluation metrics and numerical results in Section 4.
3.5 Model Performance
@@
4 Results
4.1 VLM prompting pipeline performance
The original dataset included 578 building facades from Visit 1 and 621 from Visit 2. After filtering out images labeled as “Unknown” due to obstruction, blur, poor viewing angle, or no building present, the usable sample comprised 471 facades for Visit 1 and 443 for Visit 2. These filtered counts define the set of parcels used for model evaluation across both strategies. Model performance is compared to ground truth in several ways. The Ground Truth, predictions by the Two-stage strategy, and by the One-stage strategy are shown in Figure 4. The figure shows the occupancy status for each parcel at both visits. Ground truth vacancies declined from 85 to 77 (8-8; $9.4\%$), indicating a modest decrease in vacancies between the two field surveys. The Two-stage strategy shows the same trend but with greater magnitude (75 to 63; -12; $16.0\%$), while the One-stage strategy shows an even larger drop (70 to 54; 16-16; $22.9\%$). However, both models overestimate occupancy relative to ground truth at each visit (Two-stage: 10-10 V1 / 14-14 V2; One-stage: 15-15 V1 / 23-23 V2), with larger overestimates at V2 consistent with under-predicted vacancy. Total counted parcels differ between visits because parcels with Uncertain or missing labels are excluded, but the relative shifts still show an occupancy contraction overstated by the One-stage baseline. These distributional biases signal the net recovery bias and change agreement results reported below.
@@ 4.1 VLM prompting pipeline performance
For each strategy, we compute per-visit confusion matrices (shown in Figure 5; Not Occupied is treated as positive class) and report accuracy, precision, recall, F1, and Cohen’s $\kappa$ in Table 5. Table 5 summarizes performance with 95% CIs, and Table 4 reports paired significance tests.
Table 4: Statistical significance tests for method comparisons. McNemar tests compare per-visit error patterns; paired bootstrap tests compare metric differences.
@@
Table 5: Method performance comparison with bootstrap confidence intervals. Not Occupied is treated as the positive class.
@@
As shown in Table 5, the two-stage strategy exhibits a modest precision-recall trade-off relative to the one-stage baseline. It achieves higher recall at the cost of slightly lower precision, resulting in numerically higher F1 and Cohen’s $\kappa$ scores. However, as shown in Table 4, these differences are not statistically significant. Both McNemar’s tests on per-visit error patterns and paired bootstrap tests on the performance metrics yield non-significant p-values, with confidence intervals for the differences spanning zero. This suggests that while the two-stage method yields slightly higher recall and agreement, the two strategies are statistically indistinguishable at the current operating point.
Figure 6: Per-visit overall accuracy: (a) One-stage strategy (95.5% V1, 92.1% V2); (b) Two-stage strategy (96.2% V1, 93.2% V2).
Overall accuracies (Figure 6) decline modestly from V1 to V2 for both strategies (Two-stage -3.0 pp (percentage point) vs One-stage -3.4 pp), with the Two-stage maintaining a consistent  0.7–1.1 pp advantage each visit. The parallel drop suggests a shared increase in classification difficulty (e.g., more subtle post-repair cues) rather than strategy-specific drift.
Direct comparison of the two strategies focuses on expected advantages of the two-stage (vision and reasoning) method: fewer false Not Occupied labels at a single visit (improved Occupied recall), higher precision for Deteriorated changes by propagating fewer spurious vacancies, improved precision for Recovered (fewer artificial recoveries caused by an incorrect Not Occupied baseline at V1), and higher overall change $\kappa$. Disagreements between strategies are flagged for manual review and spatially highlighted.
@@
4.2 Occupancy change analysis
We analyze parcel-level occupancy change between V1 and V2 using the four classes defined in Table 3 (Stable Occupied, Stable Not Occupied, Recovered, Deteriorated). Figure 9 (a)–(c) show Recovered / Deteriorated counts: ground truth 20 / 6 (net +14), Two-stage 23 / 9 (net +14; matches net recovery), One-stage 24 / 6 (net +18; overstates recovery by +4). Thus, both strategies capture the direction of net improvement; the One-stage inflates recoveries, while the Two-stage balances additional true recoveries with extra deteriorations.
@@ 4.2 Occupancy change analysis
Figure 9: Change analysis. (a) Ground-truth Recovered vs. Deteriorated counts (n=26). (b) One-stage predicted changes (n=30). (c) Two-stage predicted changes (n=32).
Table 6: Parcel-level occupancy change agreement categories (confident parcels with both visits).
@@
GT and Two-stage Agree 8 3.3
GT and One-stage Agree 4 1.7
@@
Spatial heterogeneity and remaining uncertainty are visualized in Figure 10, which maps (a) Visit 1 ground-truth occupancy, (b) Visit 2 ground-truth occupancy, (c) ground-truth change classes (Recovered, Deteriorated, Unchanged), and (d) spatial distribution of Two-stage prediction accuracy (both visits correct, one visit correct, both wrong). This multi-panel view highlights localized pockets of recovery and deterioration and reveals where residual model errors cluster geographically, indicating priority areas for data or prompt refinement.
Figure 10: Geographic distribution of occupancy and model accuracy: (a) Visit 1 ground-truth status; (b) Visit 2 ground-truth status; (c) ground-truth change classes (Recovered, Deteriorated, Unchanged); (d) Two-stage prediction accuracy by parcel (both visits correct, one visit correct, both wrong).
@@
We assessed spatial autocorrelation with the global Moran’s $I$ for four spatial fields: parcel occupancy at Visit 1 and Visit 2, the ground-truth change field, and a binary accuracy indicator (1 = correct, 0 = incorrect). Positive $I$ with small $p$-values indicates spatial clustering rather than spatial randomness. Table 7 summarizes the results.
Table 7: Global Moran’s II statistics for parcel-level layers. Positive values indicate spatial clustering. The expected value under the randomization null is E[I]=1/(n1)E[I]=-1/(n-1); here $E[I]\approx-0.0042$ because $n=240$.
@@
Visit 1 occupancy status 0.3704 0.0042-0.0042 0.0000
Visit 2 occupancy status 0.2834 0.0042-0.0042 0.0000
@@
5 Discussion
Our results suggest that separating perception (attribute extraction) from decision-making (conservative reasoning) is a pragmatic design choice for post-disaster occupancy inference. While the two-stage strategy yields higher recall and slightly higher F1/Cohen’s $\kappa$ at the selected operating point, paired tests do not detect a statistically significant difference compared to the one-stage baseline. We therefore interpret the primary advantage of the two-stage design not as raw performance, but as controllability and auditability. By exposing intermediate attributes, this approach allows operators to diagnose failure modes, tune decision policies (e.g., risk thresholds), and prioritize targeted human review in spatially clustered error pockets.
Beyond model architecture, the pipeline offers significant operational leverage. The separation of perception and reasoning mirrors the practical need to scale evidence collection: 360° video capture allowed for the rapid assessment of 914 residential parcels with limited field time. However, the performance drop in Visit 2 highlights a persistent challenge: post-repair scenes exhibit subtler cues that increase ambiguity. In such regimes, a conservative reasoning stage helps dampen the impact of uncertain visual evidence, though performance ultimately remains bounded by the quality of the underlying imagery.
6 Conclusion
Automated occupancy monitoring is crucial for tracking community recovery after disasters. We presented a street-level pipeline that pairs panoramic rectification with vision-language prompting to assess occupancy and track changes at the parcel scale. By fusing 5.6K rectified video frames with a VLM-based reasoning engine, we demonstrated a workflow that produces decision-ready, auditable outputs while successfully reproducing ground-truth net recovery trends.
We compared a transparent one-stage rule against a two-stage approach that separates perception from reasoning. While both strategies achieved high agreement with ground truth, the two-stage design offers superior interpretability. It allows for the isolation of visual perception errors from decision logic, facilitating the ”tuning” of conservatism required in sensitive post-disaster contexts. Crucially, our spatial analysis revealed that residual errors are not random but spatially clustered; this finding validates the use of this pipeline to direct human QA/QC efforts to specific neighborhoods rather than diffuse random audits.
Limitations remain regarding geographic generalization and environmental variability, motivating future work in multi-region validation and temporal transition modeling. Ultimately, this framework provides a scalable alternative to door-to-door inspections, shifting field effort toward rapid right-of-way capture and providing recovery managers with an auditable, spatially-aware shortlist of recovery progress.